Light of intensity I0 and polarized parallel to the transmission axis of a polarizer is incident on an analyzer.

(a) If the transmission axis of the analyzer makes an angle of 42
Answer #1


a)

I = (Io/2)*cos42^2

I = 0.27 Io


b) I / Io = costheta^2 = 7/8

theta = 20.7